# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v1

@@ 4.2 Nature of Data
Format: Network monitoring is primarily based on collecting three types of data: packets, network flows, and logs [63]. Accordingly, NID datasets are typically distributed these formats. Some also provide auxiliary data like end-host logs, routing tables, or SNMP databases. This survey outlines the four most common data formats and includes the category of other for specific cases, described as follows:
@@
  • flows: A flow is a sequence of packets with some common properties that pass through a network device [65]. These properties are typically considered as a 5-tuple: source IP, destination IP, source port, destination port, and a transport protocol identifier, used for packet aggregation. Aggregation can either be uni-directional (flows (uni)) or bi-directional (flows (bi)), capturing both sides of the connection in a single flow entry. Collected traffic features depend on flow exporters [66], such as general-purpose tools like YAF [67], nProbe [42], NFStream [68], or security-specific tools such as Argus [69], Zeek [70], and CICFlowMeter [39] as well as flow protocols they use (primarily NetFlow v5/v9 [65], IPFIX [71], sFlow [72], or OpenFlow [73]). Essential flow features (Table 3) can be found in most flow-based datasets. However, they are typically insufficient for cybersecurity, leading to the usage of extended versions with additional statistics like NetFlow v9 [74] or CICFlowMeter [75] feature sets. Most frequently, these datasets are distributed in a textual comma-separated-values (CSV) format.
@@
In general, packet-based datasets offer more detailed information because they provide exact contents and timestamps of every packet in contrast to flow aggregates. If packets’ payloads are not stripped, they also offer a possibility for deep packet inspection (DPI) [77], a technique for full packet analysis. However, with rising traffic volumes and widespread encryption, flow-based NIDSs have gained popularity for being unaffected by encryption and offering higher traffic throughput [78, 79].
@@ 4.3 Data Volume
This category analyzes the volume of the data, which can be examined from two perspectives: how many data points (i.e., raw packets or feature vectors) the dataset has and how long its capture was performed.
Although this information is typically declared in datasets’ documentation, we did not rely on it and rather downloaded and analyzed the data ourselves due to occasional inconsistencies between data and its documentation. For datasets with pre-defined training and testing splits, we report these volumes separately using a ‘+’ sign. The properties are defined as follows:
Size: Specifies the number of packets (pkts), flows (flows), or records (recs) in the dataset. For their representation, we use metric suffixes like kilo (k) = $10^{6}$ and mega (M) = $10^{9}$. A comma (“,”) represents a delimiter between thousands. In order to keep the table concise, we omit listing the size of associated log files, as the survey is focused on network data.
Duration: This property represents the dataset capture duration. While dataset documentation often defines duration as the time span between the first and last timestamps, this approach may be misleading due to the presence of potential temporal gaps. Therefore, this survey aims to report the actual data duration by disregarding such gaps.
Supposing the dataset entries are sorted chronologically, with $t_{i}$ as a timestamp of record $x_{i}$ and a difference between two adjacent records as $\tau_{i}=t_{i}-t_{i-1}$, a temporal gap in the data occurs if $\tau_{i}>\epsilon$. If no temporal gap exists ($\forall i:\tau_{i}\leq\epsilon$), i.e., the network activity is present throughout the entire dataset time span, we refer to the dataset as continuous (cont.). On the other hand, a dataset is discontinuous (discont.) if at least one capture gap is present ($\exists i:\tau_{i}>\epsilon$). In this matter, a capture duration $D$ (Equation 1) is calculated by ignoring gaps longer than $\epsilon$.
@@
In our survey, data continuity is based on $\epsilon$ relative to the dataset span (the first and last timestamps difference), as, for instance, a 10-minute gap is more tolerable in a month-long dataset than in a one-hour dataset. Therefore, we set $\epsilon$ as 1% of the analyzed dataset’s absolute time span. Additionally, assessing continuity for datasets with a pre-defined train-test split is performed separately with different $\epsilon$ for each split, as merging them would be semantically incorrect.
@@
In addition to determining how the data duration is computed, information about continuity also aids in assessing whether the data would be suitable for systems relying on time dependencies between samples. If the data was discontinuous, these systems could not utilize their dependency-analysis mechanism, likely resulting in biased evaluation results.
@@ 4.4 Network Properties
The General information category has distinguished the character of environment as real-world, emulated, or synthetic for both types of traffic. This category provides additional information about the environment – the underlying network type and whether or not the capture covers the whole network.
Network type: This property defines the network’s size and type to indicate expected traffic patterns. We specify three network sizes: 1) small(up to 99 hosts) – a small office/home office (SOHO), small company, or a household network, 2) medium(100-499 hosts), and 3) large(500+ hosts) – a large enterprise or an internet service provider (ISP). In addition to size, we also report the network’s type, with the most common being academic, enterprise, industrial, military, wireless, cloud, honeypot, internet of things (IoT), software-defined network (SDN), peer-to-peer (P2P), and internet service provider backbone (ISP). This information is omitted if no specific type is provided in the documentation. Datasets with traces combined from multiple networks are assigned the value of “mixed”.
Complete capture: Specifies whether the dataset contains data from the entire underlying network. In general, it is desirable to capture all network communication to ensure traffic diversity and realism. However, some datasets capture traffic only from a single host (e.g., honeypot) or a specific set of hosts due to privacy or practical reasons. Such a setup might skew the traffic patterns (e.g., honeypots would receive significantly more malicious traffic than regular hosts [82]). Therefore, specialized evaluation processes or NIDS types should be employed when dealing with incomplete captures.
@@ 4.5 Evaluation
Split: Defines whether the dataset offers pre-split train and test subsets for evaluation. This is desirable, as it allows transparent comparisons of different NIDSs evaluated on the same data. On the other hand, datasets without a train-test split have to be manually split by their users, which can lead to evaluation bias and selective reporting [14, 83].
Labels: Specifies whether the dataset is labeled – a crucial property in analyzing and comparing NIDSs. We assign the value yes only if an unambiguous mapping between instances and classes exists. This applies to direct labels in feature vectors but also to cases when classes can be separated via filters (e.g., IP-based packet labeling in 2017-SUEE8 [84]) or directory structures (e.g., CAIDA DDoS 2007 [85]).
@@ 5 Network Intrusion Datasets
This section, the core of our survey, lists and comparatively analyzes datasets for network intrusion detection. As outlined, the datasets were selected in accordance with the systematic literature review methodology discussed in Section 2 and A.1. Afterward, 13 properties defined in Section 4 were extracted through documentation review and data exploratory analysis. The analysis results – Jupyter Notebooks and scripts, are available at our GitHub repository (linked in Section 1).
The survey is summarized in Table LABEL:tab:data_survey, with each row representing a single dataset. In rare cases like NF-UQ-NIDS [40], two related datasets are listed as one entry divided by a ‘/’ symbol. The entries are sorted in ascending order by year and then alphabetically. Datasets spanning multiple years are ordered based on their last capture year but always appear before datasets captured purely in a given year. Dataset download links (functional in 2025) can be found in their associated reference.
The following subsections categorize and analyze the surveyed datasets in more detail. In order to group the data, we developed a taxonomy (Figure 3) classifying datasets as general-purpose (Section 5.1) or special-purpose (Section 5.2) based on their focus and intended use cases. Afterward, Section 5.3 outlines sources for other datasets not included in this survey. In Section 5.4, we analyze the popularity of datasets in contemporary state-of-the-art NIDS research, and finally, Section 5.5 highlights observed trends in network intrusion data research.
@@ 5.1 General-Purpose Datasets
General-purpose datasets, labeled as General, identify data for benchmarking intrusion detection systems for traditional network environments such as local area networks (LANs) or enterprise networks, offering a broader set of intrusions rather than focusing on specific attacks.
Although general-purpose datasets cannot be precisely divided into subcategories without ambiguities (Figure 3), we split this subsection into five parts. Initially, we describe the first-generation datasets produced by DARPA. Afterward, the paper looks at CIC datasets, which mostly replaced DARPA in recent research. Other general-purpose datasets available in the domain are also examined: NetFlow, newly released, as well as already established yet less commonly utilized.
@@ 5.1.1 DARPA & KDD
Datasets published by the Defense Advanced Research Projects Agency (DARPA) are considered the first generation of public IDS benchmarks. The first dataset, DARPA 1998 [81], emulated an Air Force base with thousands of hosts, includes a robust benign traffic profile along with 32 different attacks, categorized into Denial of Service (DoS), Probing, User-to-Root (U2R), and Remote-to-Local (R2L) classes. Despite having different goals (escalating privileges vs. gaining an access to a local account from a remote machine), both U2R and R2L classes are mainly performed via the same techniques: brute-force and vulnerability exploitation. Its data collection lasted nine weeks, each consisting only of working days, with a train-test split after the seventh week.
Since DARPA 1998 provides only raw packet and host log data, additional feature extraction and labeling steps are required to utilize it. In order to simplify its usage, KDD Cup 1999 (KDD'99) [87] introduced 41 pre-extracted features with a label by processing DARPA 1998 raw data. Its practicality, respected organization backing, and general lack of alternatives made it a de facto standard for IDS benchmarks for years to come. Its two derivatives, NSL-KDD [88] and GureKDDcup [86], further improve data quality by removing duplicate records and decreasing the dataset size, and enriching records with packet payload data (GureKDDcup).
Other DARPA-family datasets – DARPA 1999 [89], extending the DARPA 1998, and LLDOS (DARPA 2000) [90], focusing on the infiltration and DoS attacks using the DARPA 1999 environment, were also published soon after. Nevertheless, due to providing the data only in raw format, these datasets never gained significant attention from the community.
@@ 5.1.2 Canadian Institute for Cybersecurity
Building on the concept of profiles, Sharafaldin et al. introduced the B-profile and M-profile systems [163] and published the dataset CIC-IDS2017 [106] spanning 40 hours over 5 days. The dataset has introduced encrypted benign traffic and included new and popular intrusions like Heartbleed, SQL injection, or ARES Botnet. A similar scenario was run a year later on a virtualized AWS environment with hundreds of hosts, resulting in the CSE-CIC-IDS2018 [106] dataset. A characteristic property of both datasets was the employment of CICFlowMeter [39] for feature extraction, introducing a novel feature set for network intrusion detection of approximately 80 (the exact number depends on the tool version) features suitable for machine learning algorithms. The datasets and the feature set itself have gained much attention in the community.
@@ 5.1.3 NetFlow Datasets
The following paragraphs cover general-purpose datasets utilizing features solely from the NetFlow specification. Although they can also be categorized under the next Subsections (5.1.45.1.5), we discuss them separately due to their real-world practicality. In contrast to most NID datasets with custom features (unavailable in production networks without additional tools), NetFlow datasets rely on standardized features of the NetFlow architecture already deployed in many production networks for monitoring purposes. Therefore, NetFlow-based NIDSs are considered more practical due to lower implementation costs and more realistic benchmark results.
@@
Instead of obtaining real traffic from honeypots, UGR-16 [104] captures NetFlow traces from a Tier-3 ISP network over three months. The dataset aims to provide periodic network activity patterns such as day-and-night cycles or work days and weekend cycles. It includes only flow-detectable attacks such as DoS, port scans, or botnet traffic, either emulated or injected from other captures. Since the capture contains real traffic, its authors used signature- and anomaly-based methods to detect and label malicious traffic, identifying attacks such as UDP scanning, SSH scanning, or spam.
The works by Sarhan et al. [41, 40] have identified the issue of lacking data with real-world features and provided practical datasets based on the NetFlow protocol to address it. For this purpose, popular datasets – UNSW-NB15, Bot-IoT, ToN-IoT, and CSE-CIC-IDS2018, were re-generated via nProbe [42], providing basic (15) and extended (45) feature sets. While available standalone, the authors have also merged them, resulting in NF-UQ-NIDS and NF-UQ-NIDS-v2 [40] datasets.
@@ 5.1.4 Established Datasets
One of the most well-known datasets for NIDS research is UNSW-NB15 [99] (2015), emulated via the PerfectStorm platform [^3] PerfectStorm: https://www.keysight.com/br/pt/products/network-test/network-test-hardware/perfectstorm.html. It provides nine intrusion classes, including DoS, backdoors, reconnaissance, and worms. The traffic was captured as raw PCAP, host logs, and 49 flow-based features. However, its documentation lacks details on data generation, causing ambiguities in its interpretation. For instance, there is no description of emulated vulnerabilities, types of performed DoS attacks, or their parameters. For this reason, conclusions drawn by evaluations using this dataset might be less sound. PerfectStorm was also utilized to generate the NGIDS-DS [103] dataset, which also shares the issues related to limited documentation. In addition, it provides only raw data labeled indirectly via host logs, further degrading its usability.
@@ 5.1.5 New Datasets
In recent years, several new general-purpose datasets have been released. Due to their novelty, they are expected to offer the newest benign traffic patterns and current intrusions. Therefore, the results obtained using them as benchmarks should be closest to reality. Since they were released only recently, the community has not yet thoroughly analyzed them and “decided” on their popularity, as they are not yet commonly referenced in recent scientific publications. For this reason, a separate paragraph is provided with an introduction and a brief description of each newly-released dataset.
@@
OD-IDS2022 [151] (2022) offers a diverse emulation of modern attacks (28) captured in a virtualized network. The environment contains external and internal attacker networks targeting an Ubuntu Apache HTTP server. The features were extracted using CICFlowMeter. However, details of the environment and normal traffic profiles are not provided.
UWF-ZeekData22 [149] (2022) contains traces from a cyber wargaming course at the University of West Florida. In the course, red (offensive) and blue (defensive) teams are tasked to attack/defend the parts of the university network infrastructure. The traffic features were extracted by Zeek and labeled based on the MITRE ATT&CK [167] phases using participants’ logs. However, 99.9999% of malicious traffic involves reconnaissance and discovery, leaving too few samples for other phases. Therefore, the dataset consists mostly of scanning traffic, making it unsuitable for systems focusing on other intrusion types.
VHS-22 [148] (“Very Heterogenous Set of Network Traffic Data”, 2004 – 2022) combines five datasets – ISOT Botnet, CTU-13, Booters, CICIDS-2017, and captures from the Malware Traffic Analysis (MTA) project [168]. Their traffic was replayed, and timestamps reset on 2022-01-01, thus resembling a 23-day-long continuous capture with 115 classes. The data is provided as unidirectional flows extracted via a custom network probe. Its high diversity will likely prove challenging for some ML methods. For instance, anomaly detectors might struggle to establish a sense of a traffic normality profile.
@@
Finally, TII-SSRC-23 [154] (2023) promotes data diversity by providing eight distinct traffic types. They are divided into six benign (network use-cases like audio or video streaming, text, and background) and 26 malicious subtypes (brute-force, DoS, information gathering, and Mirai botnet). The data are provided in the raw PCAP and the bidirectional flow format produced by CICFlowMeter. However, the dataset is heavily imbalanced, with benign traffic comprising only 0.015% (1300 out of 8.7 million) flows, posing challenges for its utilization.

### v2

@@ 4.2 Nature of Data
Format: Network monitoring is primarily based on collecting three types of data: packets, network flows, and logs [64]. Accordingly, NID datasets are typically distributed in these formats. Some also offer auxiliary data like end-host logs, routing tables, or SNMP databases. This survey outlines the four most common data formats and includes the category of other for specific cases, as described below:
@@
  • flows: A flow is a sequence of packets with some common properties that pass through a network device [66]. These properties are typically considered as a 5-tuple: source IP, destination IP, source port, destination port, and a transport protocol identifier, used for packet aggregation. Aggregation can either be uni-directional (flows (uni)) or bi-directional (flows (bi)), recording both sides of the connection in a single flow entry.
    Collected traffic features depend on flow exporters [67], including general-purpose tools like YAF [68], nProbe [44], and NFStream [69], as well as security-focused tools like Argus [70], Zeek [71], and CICFlowMeter [41]. These exporters rely on various flow protocols, primarily NetFlow v5/v9 [66], IPFIX [72], sFlow [73], and OpenFlow [74]. While essential flow features (Table 3) are commonly found in most flow-based datasets, they are often insufficient for cybersecurity applications. As a result, extended versions with additional statistics, such as NetFlow v9 [75] or CICFlowMeter [76] feature sets, are commonly used. Most frequently, these datasets are distributed in a textual comma-separated-values (CSV) format.
@@
In general, packet-based datasets provide more detailed information because they provide exact contents and timestamps of every packet in contrast to flow aggregates. If packet payloads are not stripped, they also enable deep packet inspection (DPI) [78], a technique for full packet analysis. However, with increasing traffic volumes and widespread encryption, flow-based NIDSs have gained popularity for being unaffected by encryption and offering higher traffic throughput [79, 80].
@@ 4.3 Data Volume
This category evaluates the volume of the data from two perspectives: the number of data points (i.e., raw packets or feature vectors) and the data capture duration.
Although this information is typically declared in dataset documentation, we did not rely on it. Instead, we downloaded and analyzed the data ourselves due to occasional inconsistencies between the data and its documentation. For datasets with predefined training and testing splits, we report these volumes separately using a “+” sign. The properties are defined as follows:
Size: Specifies the number of packets (pkts), flows (flows), or records (recs) in the dataset. For their representation, we use metric suffixes like kilo (k) = $10^{6}$ and mega (M) = $10^{9}$. A comma sign (“,”) represents a delimiter between thousands. In order to keep the table concise, we omit listing the size of associated log files, as the survey is focused on network data.
Duration: This property represents the dataset capture duration. While dataset documentation often defines duration as the time span between the first and last data timestamps, this approach may be misleading due to the presence of potential temporal gaps. Therefore, this survey aims to report the actual data duration by disregarding such gaps.
Supposing the dataset entries are sorted chronologically, with $t_{i}$ as a timestamp of record $x_{i}$ and a difference between two adjacent records as $\tau_{i}=t_{i}-t_{i-1}$, a temporal gap in the data occurs if $\tau_{i}>\epsilon$. If no temporal gap exists ($\forall i:\tau_{i}\leq\epsilon$), i.e., the network activity is present throughout the entire dataset time span, we refer to the dataset as continuous (cont.). Otherwise, a dataset is discontinuous (discont.) if at least one capture gap is present ($\exists i:\tau_{i}>\epsilon$). In this matter, an actual capture duration $D$ (Equation 1) is calculated by ignoring gaps longer than $\epsilon$.
@@
In this survey, data continuity is based on $\epsilon$ relative to the dataset span (the first and last timestamps difference), as, for instance, a 10-minute gap is more acceptable in a month-long dataset than in a one-hour dataset. Therefore, we set $\epsilon$ as 1% of the analyzed dataset’s absolute time span. Additionally, assessing continuity for datasets with a pre-defined train-test split is performed separately with different $\epsilon$ for each split, as merging them would be semantically incorrect.
@@
In addition to determining how the data duration is computed, information about continuity also aids in assessing whether the data would be suitable for detection systems relying on time dependencies between samples. If the data was discontinuous, these systems could not utilize their dependency-analysis mechanism, likely resulting in biased evaluation results.
@@ 4.4 Network Properties
The General information category distinguished the character of environment as real-world, emulated, or synthetic for both types of traffic. In this category, we provide additional information about the environment – the underlying network type and whether the capture covers the entire network.
Network type: This property defines the network’s size and type to indicate expected traffic patterns. We categorize network sizes into three groups: 1) small(up to 99 hosts) – a small office/home office (SOHO), small company, or a household network, 2) medium(100-499 hosts), and 3) large(500+ hosts) – a large enterprise or an internet service provider (ISP). In addition to size, we also report the network’s type, with the most common being academic, enterprise, industrial, military, wireless, cloud, honeypot, internet of things (IoT), software-defined network (SDN), peer-to-peer (P2P), and internet service provider backbone (ISP). This information is omitted if no specific type is provided in the documentation. Datasets with traces combined from multiple networks are assigned the value of “mixed”.
Complete capture: Specifies whether the dataset contains data from the entire underlying network. In general, it is desirable to capture all network communication to ensure traffic diversity and realism. However, some datasets capture only from a single host (e.g., honeypot) or a specific set of hosts due to privacy or practical reasons. Such a setup might skew the traffic patterns (e.g., honeypots would receive significantly more malicious traffic than regular hosts [83]). Therefore, specialized evaluation processes or NIDS types should be employed when dealing with incomplete captures.
@@ 4.5 Evaluation
Split: Defines whether the dataset offers pre-split train and test subsets for evaluation. This is desirable, as it allows transparent comparisons of different NIDSs evaluated on the same data. On the other hand, datasets without a train-test split have to be manually split by their users, potentially leading to evaluation bias and selective reporting [15, 84].
Labels: Specifies whether the dataset is labeled – a crucial property in analyzing and comparing NIDSs. We assign the value yes only if an unambiguous mapping between data instances and classes exists. This applies to direct labels in feature vectors but also to cases when classes can be separated via filters (e.g., IP-based packet labeling in 2017-SUEE8 [85]) or directory structures (e.g., CAIDA DDoS 2007 [86]).
@@ 5 Network Intrusion Datasets
This section, the core of this survey, lists and comparatively analyzes datasets for network intrusion detection. As outlined, the datasets were selected in accordance with the systematic literature review methodology discussed in Section 2 and A.1. Afterward, 13 properties defined in Section 4 were extracted through documentation review and data exploratory analysis. The analysis results – Jupyter Notebooks and scripts, are available on our GitHub repository (linked in Section 1).
The survey results are summarized in Table LABEL:tab:data_survey, with each row representing a single dataset. In rare cases like NF-UQ-NIDS [42], two related datasets are listed as one entry divided by a “/” symbol. The entries are sorted in ascending order by year and then alphabetically. Datasets spanning multiple years are ordered based on their last capture year but always appear before datasets captured in a single year. Dataset download links (functional in early 2025) are in their associated references.
The following subsections categorize and analyze the surveyed datasets in more detail. In order to group the data, we developed a taxonomy (Figure 3) classifying datasets as general-purpose (Section 5.1) or special-purpose (Section 5.2) based on their focus and intended use cases. Afterward, Section 5.3 outlines sources for other datasets not included in this survey. In Section 5.4, we analyze the popularity of datasets in contemporary state-of-the-art NIDS research. Finally, Section 5.5 highlights observed trends in network intrusion data research.
@@ 5.1 General-Purpose Datasets
General-purpose datasets, labeled as General, provide data for benchmarking intrusion detection systems for traditional network environments such as local area networks (LANs) or enterprise networks. They offer a more diverse set of intrusions rather than focusing on specific attacks.
Although general-purpose datasets cannot be precisely divided into subcategories without ambiguity (Figure 3), we organize this subsection into five parts. Initially, we describe the first-generation datasets produced by DARPA. Next, we examine CIC datasets, which have largely replaced DARPA in recent research. Finally, we review other general-purpose datasets, including NetFlow, newly released, as well as already established but less commonly utilized.
@@ 5.1.1 DARPA & KDD
Datasets published by the Defense Advanced Research Projects Agency (DARPA) are considered the first generation of public IDS benchmarks. The first dataset, DARPA 1998 [82], which emulated an Air Force base with thousands of hosts, includes a robust benign traffic profile along with 32 different attacks, categorized into Denial of Service (DoS), Probing, User-to-Root (U2R), and Remote-to-Local (R2L) classes. Despite having different goals (escalating privileges vs. gaining an access to a local account from a remote machine), both U2R and R2L classes are mainly performed via brute-force and vulnerability exploitation. Its data collection lasted nine weeks during weekdays, with a train-test split after the seventh week.
Since DARPA 1998 provides only raw packet and host log data, additional feature extraction and labeling steps are required to utilize it. In order to simplify its usage, KDD Cup 1999 (KDD'99) [88] introduced 41 pre-extracted features with a label by processing DARPA 1998 raw data. Its practicality, respected organization backing, and general lack of alternatives made it a de facto standard for IDS benchmarks for years to come. Its two derivatives, NSL-KDD [89] and GureKDDcup [87], further improve data quality by removing duplicate records and reducing the dataset size (NSL-KDD) or enriching records with packet payload data (GureKDDcup).
Other DARPA-family datasets – DARPA 1999 [90], extending the DARPA 1998, and LLDOS (DARPA 2000) [91], focusing on the infiltration and DoS attacks using the DARPA 1999 environment, were also published soon after. Nevertheless, because these datasets are provided only in raw format, they have never never gained significant attention from the community.
@@ 5.1.2 Canadian Institute for Cybersecurity
Building on the concept of profiles, Sharafaldin et al. introduced the B-profile and M-profile systems [164] and published the dataset CIC-IDS2017 [107], spanning 40 hours over five days. The dataset introduced encrypted benign traffic and new widespread intrusions like Heartbleed, SQL injection, or ARES Botnet. A similar scenario was executed a year later on a virtualized AWS environment with hundreds of hosts, resulting in the CSE-CIC-IDS2018 [107] dataset. A characteristic property of both datasets was the employment of CICFlowMeter [41] for feature extraction, introducing a novel feature set for network intrusion detection of approximately 80 (the exact number depends on the tool version) features suitable for machine learning algorithms. The datasets and the feature set itself have gained much attention in the community.
@@ 5.1.3 NetFlow Datasets
The following paragraphs cover general-purpose datasets utilizing features exclusively from the NetFlow specification. While they could also be categorized under the next subsections (5.1.45.1.5), we discuss them separately due to their real-world relevance. In contrast to most NID datasets with custom features (unavailable in production networks without additional tools), NetFlow datasets rely on standardized features of the NetFlow architecture already deployed in many networks for monitoring purposes. Therefore, NetFlow-based NIDSs are considered more practical due to lower implementation costs and more realistic benchmark results.
@@
Instead of obtaining real traffic from honeypots, UGR-16 [105] captures NetFlow traces from a Tier-3 ISP network over three months. The dataset aims to provide periodic network activity patterns such as day-and-night cycles or weekdays and weekend cycles. It includes only flow-detectable attacks such as DoS, port scans, or botnet traffic, either emulated or injected from other captures. Since the capture contains real traffic, its authors used signature- and anomaly-based methods to detect and label malicious traffic, identifying attacks such as UDP scanning, SSH scanning, or spam.
The works by Sarhan et al. [43, 42] have identified the issue of lacking data with real-world features and provided practical datasets based on the NetFlow protocol to address it. For this purpose, popular datasets – UNSW-NB15, Bot-IoT, ToN-IoT, and CSE-CIC-IDS2018 – were regenerated using nProbe [44], providing basic (15) and extended (45) feature sets. While available individually, the authors also merged them, resulting in the NF-UQ-NIDS and NF-UQ-NIDS-v2 [42] datasets.
@@ 5.1.4 Established Datasets
One of the most prominent datasets for NIDS research is UNSW-NB15 [100] (2015). Emulated using the PerfectStorm platform [^3] PerfectStorm: https://www.keysight.com/br/pt/products/network-test/network-test-hardware/perfectstorm.html, it includes nine intrusion classes, such as DoS, backdoors, reconnaissance, and worms. The data is provided as raw PCAP, host logs, and 49 flow-based features. However, its documentation lacks details on data generation, leading to ambiguities in its interpretation. For instance, there is no description of the emulated vulnerabilities, types of attacks, or their parameters. Recently, issues with mislabeling, simulation artifacts, and lack of diversity have also been discovered [8]. For this reason, conclusions drawn by evaluations using this dataset might be less sound. PerfectStorm was also used to generate NGIDS-DS [104], which also shares these documentation issues. In addition, it only provides raw data labeled indirectly via host logs, further limiting its usability.
@@ 5.1.5 New Datasets
In recent years, several new general-purpose datasets have been released. Due to their novelty, they are expected to offer the newest benign traffic patterns and current intrusions. Therefore, the results obtained using them as benchmarks should be closer to reality. Since they were released only recently, the community has not yet thoroughly analyzed them and “decided” on their popularity, as they are not yet commonly referenced in recent scientific publications. For this reason, a separate paragraph is provided with an introduction and a brief description of each newly-released dataset.
@@
OD-IDS2022 [152] (2022) offers a diverse emulation of modern attacks (28) captured in a virtualized network. The environment contains external and internal attacker networks targeting an Ubuntu Apache HTTP server. The features were extracted using CICFlowMeter. However, details of the environment and normal traffic profiles are lacking.
UWF-ZeekData22 [149] (2022) contains traces from a cyber wargaming course at the University of West Florida. During the course, red (offensive) and blue (defensive) teams were tasked to attack/defend the parts of the university network. The traffic features were extracted by Zeek and labeled based on the MITRE ATT&CK [168] phases using participants’ logs. However, 99.9999% of malicious traffic involves reconnaissance and discovery, leaving too few samples for other phases. Therefore, the dataset consists mostly of scanning traffic, making it unsuitable for systems focusing on other intrusion types.
VHS-22 [150] (“Very Heterogenous Set of Network Traffic Data”, 2004 – 2022) combines five datasets – ISOT Botnet, CTU-13, Booters, CICIDS-2017, and captures from the Malware Traffic Analysis (MTA) project [169]. Their traffic was replayed, and timestamps reset on 2022-01-01, thus resembling a single capture (including gaps) with 115 classes. The data is provided as unidirectional flows extracted via a custom network probe. Its high diversity will likely prove challenging for some ML methods, as anomaly detectors might struggle to establish a sense of a traffic normality profile.
@@
Finally, TII-SSRC-23 [155] (2023) promotes data diversity by providing eight distinct traffic types. They are divided into six benign (network use-cases like audio or video streaming, text, and background) and 26 malicious subtypes (brute-force, DoS, information gathering, and Mirai botnet). The data are provided as the raw PCAP and the bidirectional flow format produced by CICFlowMeter. However, the dataset is heavily imbalanced, with benign traffic comprising only 0.015% (1300 out of 8.7 million) flows, posing challenges for its utilization.